OpenStudy (anonymous):

lim x→∞ (e^(−3x)cos x)

OpenStudy (anonymous):

\[\large\lim_{x\to\infty}e^{-3x}\cos x=\lim_{x\to\infty}\frac{\cos x}{e^{3x}}\] Use the fact that cosine is bounded, i.e. \(-1\le\cos x\le1\).

OpenStudy (anonymous):

last one? can you help with this one? lim x→(3π/2)+ (etan x)

OpenStudy (anonymous):

Is that supposed to be \(e^{\tan x}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Using the continuity of the exponential function, you can rewrite:\[\large\lim_{x\to(3\pi/2)^+}e^{\tan x}=\large e^{\left(\displaystyle\lim_{x\to(3\pi/2)^+}\tan x\right)}\]